### v2

# Faster Model Predictive Control via Self-Supervised Initialization Learning

Abstract
Optimization for robot control tasks, spanning various methodologies, includes Model Predictive Control (MPC). However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability. Prior works in speeding up the optimization have limitations on optimizing MPC running time directly and generalizing to hold out domains. To overcome this challenge, we develop a novel framework aiming at expediting optimization processes directly. In our framework, we combine offline self-supervised learning and online fine-tuning to improve the control performance and reduce optimization time. We demonstrate the success of our method on a novel and challenging Formula 1 track driving task. Comparing to single-phase training, our approach achieves a $19.4\%$ reduction in optimization time and a $6.3\%$ improvement in tracking accuracy on zero-shot tracks.
I INTRODUCTION
Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc. However, MPC in real-world applications is challenging due to computational time bottlenecks [8, 9, 10, 11], particularly with complex systems featuring non-convex and non-differentiable cost functions and prolonged planning horizons. [12, 13]. Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
Figure 1: Overview of our proposed algorithm. The first two blocks denote the two-phase training framework. In the first phase, we collect expert MPC demonstrations and train a warm-start policy using behavior cloning to speed up MPC. In the second phase, we fine-tune this policy within an online training framework to enhance its performance and generalizability. During testing, the proposed framework is evaluated on both training tracks and challenging zero-shot tracks, as demonstrated in the third block.
Warm starting in robot control optimization involves initializing the solver with a closer-to-optimal solution, speeding up the iterative optimization process. The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations. As shown in Figure 2, a closer initial guess reduces the distance the optimization algorithm needs to traverse to converge to the optimal solution. By starting closer to the optimum, the solver can often avoid lengthy exploration of the solution space, leading to faster convergence and reduced computational effort. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
To provide MPC with a high quality initial guess, various approaches have been adopted. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [14, 15]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [10, 16]. However, this approach may lack zero-shot generalizability. As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase, thus causing time-consuming operations to searching for the initial guess within the buffer. Alternatively Klaučo et al. propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [17]. However, this method lacks precise warm-start guesses and struggles with complex control tasks involving long planning horizons and heterogeneous observations. Moreover, it’s limited to strictly convex quadratic programs, reducing its applicability to real-world problems with non-linear or non-differentiable objectives and dynamics.
More recent approaches have adopted supervised learning to boost the performance of model predictive optimization algorithms. [18, 19] adopts behavior cloning (BC) to learn the distribution of the initial guesses for the solver. However, supervised learning cannot directly learn to minimize the computation time of MPC. Instead, all supervised learning can do is to imitate the output of the solver. Besides, BC also suffers from covariance shift, which degrades model’s performance on zero-shot scenarios [20]. [21] used reinforcement learning (RL) to train a warm-start policy to help quadrotor track simple zig-zag trajectories. Yet, RL algorithms are too data hungry to learn an optimal policy from scratch, posing challenge to complex, long-horizon tasks [22]. Furthermore, [21] is still not directly learning to minimize the computation time of MPC. Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time. Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional zero-shot testing phase. The two training phase works reciprocally to each other. On the one hand, the online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimize the computation time of MPC. On the other hand, the offline BC provides online learning algorithm with a good starting point to expedite the training process. Moreover, another advantage of using a learned policy to initialize the MPC, rather than for end-to-end control, is that it preserves the original definition of the control problem. The learned policy only affects the starting point, while MPC ensures the control solution satisfies system constraints, leading to safer operation. Our key contributions are:
  1. 1.
    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods all failed.
  2. 2.
    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $19.4\%$ higher performance in optimization time and a $6.3\%$ higher performance in tracking accuracy on challenging zero-shot tracks comparing to single-phase training.
Figure 2: An illustrative demo showing how a better initial guess improves MPC optimization time.
II Preliminaries
II-A Model Predictive Control
MPC is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs over a finite horizon, H. This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior. When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function , $J$ while respecting system dynamics and constraints. Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state. The optimization process is performed again at the next state. The MPC control law is formulated as shown below in Equations (1):
[…]
subject $\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
$U=[u_{t},...,u_{t+H-1}]\in\mathcal{U}_{j}\ \text{for all}\ j=1,\dots,n_{c_{u}}$ (1c)
$X=[x_{t},...,x_{t+H}]\in\mathcal{X}_{j}\ \text{for all}\ j=1,\dots,n_{c_{x}}$ (1d)
Equation (1a) formulates the objective function of the MPC. $H$ denotes the planning horizon, $t$ is the current step, and, $x_{t+i}$ is the system state at the $t+i$ step. Given the state, $x_{t+i}$ and the control input, $u_{t+i}$, the system dynamics model predicts the next state, $x_{t+i+1}$, using the dynamics model (1b). Equation (1c) and Equation (1d) are the constraints for the control inputs and states. $n_{c_{u}}$ and $n_{c_{x}}$ are the number of constraints on control inputs and states respectively.
III Method
In this section, we discuss in detail the proposed two-phase training framework. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
Algorithm 1 Offline Training
1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect $N$
2:Initialize neural network policy $\pi_{\theta}^{\text{warm}}$
3:$t\leftarrow 0,s\leftarrow s_{0},\mathcal{D}=\emptyset$
4:while $t<N$ do
5:     $(u_{t},u_{t+1},\cdots,u_{t+H-1})\leftarrow\pi^{\text{MPC}}(s,\vec{0})$
6:     $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s,u_{t},u_{t+1},\cdots,u_{t+H-1})\}$
7:     $s\leftarrow T(s,u_{t})$
8:     $t\leftarrow t+1$
9:end while
10:Train $\pi_{\theta}^{warm}$ with Equation (2) and $\mathcal{D}$
Algorithm 2 Online Fine Tuning
1:Input: fast MPC $\pi^{\text{MPC}}_{\text{fast}}$ with planning horizon $H$ and maximum optimization iteration $N_{fast}$, environment transition $T$, pre-trained warm-start policy $\pi_{\theta}^{warm}$
2:for each RL training iteration do
3:     Perceive an observation $s$
4:     $\hat{U}_{t}=(\hat{u}_{t},\hat{u}_{t+1},...,\hat{u}_{t+H-1})\leftarrow\pi^{warm}(s)$
5:     $pos_{i+1}^{car}=M_{dynamics}((s_{i},\hat{u}_{i}))|_{i=t}^{H-1}$
6:     Compute accumulated $xte$ using Equation (5)
7:     $U_{t}^{\text{MPC}}=(u_{t},u_{t+1},...,u_{t+H-1})\leftarrow\pi^{\text{MPC}}_{\text{fast}}(s,\hat{U}_{t})$
8:     Calculate reward in RL using Equation (4)
9:     $L_{imitation}=MSE(U_{t}^{MPC},U_{t}^{guess})$
10:     Compute training loss $L$ in Equation (3)
11:     $s\leftarrow T(s,u_{t})$
12:     Update $\pi_{\theta}^{warm}$ with $L$
13:end for
III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs first. The expert MPC, $\pi^{MPC}$ aims to control the agent to finish the task without relying on a warm-started initial guess for the control action sequence. At each step, $\pi^{MPC}$ observes a state $s$ and takes an all-zero vector as the initial guess and optimizes that initial guess to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). Then the state-action pair, $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in the dataset $\mathcal{D}$ (line 6). The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8). During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to ensure it achieves a high control performance by disabling the “early stop” for the expert MPC.
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [23]. Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10).
BC is a simple yet effective algorithm to learn from demonstrations. The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$. BC learns a control policy, $\pi_{\theta}$, by minimizing the Mean Squared Error (MSE) in the demonstration set, shown in Equation (2). Our algorithm utilizes pre-collected MPC data to perform BC as a warm start for the MPC to improves the MPC’s optimization time.
$\theta^{*}=\min_{\theta}{\sum_{\tau^{i}\in\mathcal{D}}\sum_{t=0}^{T}{(\pi_{\theta}^{warm}(s^{i})_{t}-u_{t}^{i})^{2}}}$ (2)
III-B Online Fine Tuning
Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through BC. This phase combines the best aspects of RL and Dataset Aggregation (DAgger) [24] algorithm. RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow A$ , is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
DAgger builds upon BC by incorporating online interaction with the environment and online querying of the expert. Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training. This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
In our framework, DAgger and RL address the covariance shift problem caused by offline BC. Additionally, RL not only fine-tunes the warm-start solution but also has the unique advantage of directly optimizing MPC computation time. We choose to leverage Twin Delayed Deep Deterministic Policy Gradient (TD3) [25] as the RL algorithm to fine-tune our pre-trained trajectory prediction model. The trajectory prediction model acts as a policy network in the TD3 framework. DAgger is integrated as a term in actor loss in the RL training, as shown in Equation (3). $B$ is the batch size and $Q_{\theta}$ is the critic network. $L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. Note that DAgger is only added when the quality of the initial guess is lower than a certain threshold. Because we only need to query the expert MPC when the initial guess is of low quality and calling the expert MPC at each step could be time-consuming. $\lambda$ is the weight coefficient between the RL loss and imitation loss.
$L$ $=\lambda\cdot L_{actor}+(1-\lambda)\cdot L_{imitation}$ (3)
$L_{\text{actor}}$ $=-\frac{1}{B}\sum_{i=1}^{B}Q_{\theta}(s_{i},\pi^{\text{warm}}(s_{i}))\cdot L_{value}$
In our driving task, the reward at each step of the RL training is shown in Equation (4). The first term is the negative MPC optimization time, and the second term is the negative of the Cross Track Error ($xte$) over the planning horizon, $H$. The $xte$ is defined in Equation (5). At each step, $\pi_{\theta}^{warm}$ will output an action sequence of length $H$ based on the current state vector, $s_{t}$. Then, the dynamics model, $M_{dynamics}$, is used to calculate the future positions of the agent. The $xte$ is computed between the reference trajectory and the future positions of the vehicle, $pos_{i}^{car}$. The reward obtained at each step directly influences $L_{RL}$ by contributing to the calculation of the advantage estimate, which measures the discrepancy between the observed reward and the expected value of the state-action pair. This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (4)
$xte=distance(pos^{car}_{t},\mathbf{p}_{i}(x_{i},y_{i})^{closest})$ (5)
This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by maximizing the tracking accuracy. This helps to address the sub-optimality problem caused by BC.
IV EXPERIMENTS
In this section, we will discuss in detail the setup of the Formula 1 track domains, followed by the MPC design in this testing domain. Additionally, we will present experimental results from both training and zero-shot tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
IV-A Experiment Setup
The experiments are done in high-speed Formula 1 tracks [26] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively. The reference trajectory is represented as a set of waypoints in the center of the track. The track size is downscaled to a 10:1 ratio to make each lap a reasonable length. The reference speed for the vehicle in the expert MPC is $10m/s$. Given the downscaled size of the track data, the vehicle could actually speed up to $100m/s$ in the 1:1 scale track. The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions.
The three tracks used for demonstration collection and training are shown in Figure 3(a). The six zero-shot tracks are not presented to the vehicle before testing to demonstrate our algorithm’s zero-shot generalizability. During training, the expert MPC is first rolled out on training tracks to collect demonstrations for offline training introduced in Section III-A. Then, the warm-start policy is fine-tuned on the same three tracks using the algorithm introduced in Section III-B.
MPC is designed to control the acceleration and the steering of the vehicle to track the reference trajectory. The dynamics model of the vehicle, $M_{dynamics}$, is shown in Equation (6). $(x_{t}^{car},y_{t}^{car})$ is the global position of the vehicle at time step $t$. $v_{t}$ and $yaw_{t}$ are the speed ($m/s$) and yaw angle ($rad$) of the vehicle at time step $t$. $a_{t}$ and $\theta_{t}^{\text{steering}}$ are acceleration ($m/s^{2}$) and steering angle ($rad$) inputs to the vehicle at time step $t$. $L$ is the wheelbase of the vehicle. Here we use Ford Mustang’s dimensions, $L=2.89m$. And $dt=0.02$ is the length of each time step which represents the duration between successive updates of control inputs and system states.
$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt$ (6a)
$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt$ (6b)
$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt$ (6c)
$v_{t+1}=v_{t}+a_{t}\cdot dt$ (6d)
The MPC objective function is composed of five parts as shown in Equation (7). The first two terms are $xte$ and Error in Heading ($eth$). $xte$ is computed using Equation 5. $eth$ is computed using Equation (8), which denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. $v_{t}^{\text{ref}}$ is the desired speed of the vehicle. The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead. Since the calculation of $xte$ involves finding the closest point on a reference trajectory to the car’s current position, it introduces discontinuities and non-differentiability into the MPC objective function. As a result, only gradient-free, sampling-based solvers are suitable for this problem [27, 28]. Here we choose to use the gradient-free solver COBYLA [29] in our experiments.
Furthermore, given the high speed of the vehicle and the presence of multiple sharp turns on the track, gradient-free MPC solvers without a good initial guess of the control solution will struggle to optimize the trajectory in real time, making real-world implementation challenging. And we show in Section IV-C gradient-free MPC solver only succeeds in simple IMS track and fails at all of the complex tracks.
$J_{t}$ $=\sum_{i=t}^{t+H-1}(w_{0}\cdot xte_{i}^{2}+w_{1}\cdot eth_{i}^{2}$ (7)
$\quad+w_{2}\cdot(v_{i}-v_{t}^{\text{ref}})^{2}$
$\quad+w_{3}\cdot(steer_{i}-steer_{i-1})^{2}$
$\quad+w_{4}\cdot(throttle_{i}-throttle_{i-1})^{2})$
$eth=abs\left(yaw^{car}_{t}-\arctan\left({\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}}}\right)\right)$ (8)
We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$ to avoid MPC wasting time on optimizing a trajectory that is already good enough. Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.05 seconds in worst-case scenarios. As, 20Hz running rate of MPC is validated by [30] for vehicle path tracking.
(a) Training tracks
(b) Zero-shot tracks (complex)
(c) Zero-shot tracks (simple)
Figure 3: Tracks tested during training and testing.
IV-B Baselines and Metrics
During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step. We compare our algorithm’s performance against five types of initial guess policies: All Zero: all-zero initial guesses, Previous Solution: initial guesses derived from the MPC solution at the previous step, Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training, Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC, and Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger. To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
IV-C Experiment Results
We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)). The results are shown in Figure 4 and Figure 5, respectively. On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy. Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely through offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality covariance shift problem caused by BC.
[…]
The results also reveal that initialize an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane when facing any sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
Figure 4: Experiment results on training tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
Figure 5: Experiment results on the six zero-shot testing tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
Furthermore, the Ours w/o RL algorithm also fails on both training and zero-shot domains. This is because the absence of offline BC requires the algorithm to explore the solution space from scratch, causing the online learning algorithm to take significantly longer to converge to a satisfactory solution. Given the same amount of training time, the quality of the policy learned by Ours w/o RL lags behind that of the policy learned through the two-phase training algorithm. This demonstrates the effectiveness of offline BC in accelerating online training.
To further inspect the relative performance of our approach when our baselines performed better, i.e. on simpler tracks, we conducted further testing on six warm-start policies on a simpler zero-shot track, IMS, as depicted in Figure 3(c),   similar to the path-tracking scenario in [31]. The results are summarized in Figure 6. Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
Figure 6: Experiment results on the simple IMS track.
In summary, our empirical results support that:
  • MPC with warm-start policy trained by either offline learning or a combination of offline and online fine-tuning significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
  • The offline training algorithm significantly warm-starts the online training algorithm and make it take much less training time to get a satisfying performance.
  • The warm-start policy trained with our two-phase learning framework reduces MPC optimization time and $xte$ on both training and zero-shot tracks, showing the capability of our online fine-tuning algorithm in addressing the suboptimality problem.
  • The two-phase training algorithm outperforms in zero-shot scenarios, demonstrating superior generalizability.
V LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are still limitations and avenues for future research that merit consideration. Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces. Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI. Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [32]. MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost. Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution. This reduces the iteration count needed for optimization, aligning with early stopping criteria. While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
VI CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver. Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and improving path tracking accuracy. This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.