### v3

# Faster Model Predictive Control via Self-Supervised Initialization Learning

Abstract
Model Predictive Control (MPC) is widely used in robot control by optimizing a sequence of control outputs over a finite-horizon. Computational approaches for MPC include deterministic methods (e.g., iLQR and COBYLA), as well as sampling-based methods (e.g., MPPI and CEM). However, complex system dynamics and non-convex or non-differentiable cost terms often lead to prohibitive optimization times that limit real-world deployment. Prior efforts to accelerate MPC have limitations on: (i) reusing previous solutions fails under sharp state changes and (ii) pure imitation learning does not target compute efficiency directly and suffers from suboptimality in the training data. To address these, We propose a warm-start framework that learns a policy to generate high-quality initial guesses for MPC solver. The policy is first trained via behavior cloning from expert MPC rollouts and then fine-tuned online with reinforcement learning to directly minimize MPC optimization time. We empirically validate that our approach improves both deterministic and sampling-based MPC methods, achieving up to 21.6% faster optimization and 34.1% more tracking accuracy for deterministic MPC in Formula 1 track path-tracking domain, and improving safety by 100%, path efficiency by 12.8%, and steering smoothness by 7.2% for sampling-based MPC in obstacle-rich navigation domain. These results demonstrate that our framework not only accelerates MPC but also improves overall control performance. Furthermore, it can be applied to a broader range of control algorithms that benefit from good initial guesses.
I INTRODUCTION
Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc. While gradient-based MPC is highly effective when the dynamics and cost functions are smooth and differentiable, applying MPC in real-world settings remains challenging. Computational bottlenecks often arise in systems with non-convex or non-differentiable costs and long planning horizons, where gradient-free solvers are typically employed [8, 9, 10, 11].
Figure 1: Overview of our proposed algorithm. The first two blocks denote the two-phase training framework. In the first phase, we collect expert MPC demonstrations and train a warm-start policy using behavior cloning to speed up MPC. In the second phase, we fine-tune this policy within an online training framework to enhance its performance and generalizability. During testing, the proposed framework is evaluated on both training tracks and challenging zero-shot tracks, as demonstrated in the third block.
Dynamic replanning in MPC requires fast computation to operate in real time, but standard MPC without warm-starting often fails to achieve sufficient speed in challenging domains [12, 13]. Warm starting addresses this by initializing the solver with a solution close to the optimum, reducing the search space and accelerating convergence. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
Prior works have employed heuristic strategies to warm-start MPC, but these methods often struggle with scalability and fail to adapt effectively to sudden changes in system states. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [12, 14]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store and leverage historical MPC solution [10, 15]. However, this approach fails to generalize to unseen states and scales.
Traditional machine learning approach has also been adopted to learn a warm-start for MPC, however, it suffers from inaccurate initialization. Klaučo et al. propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [16], but this only produces coarse guesses. These shortcomings highlight the need for a general warm-starting framework that directly learns to generate high-quality initial guesses, scales to complex domains, and improves solver efficiency under sharp dynamics and diverse conditions.
More recent approaches have attempted to predict effective warm-starts for model predictive optimization by either behavior cloning (BC) historical trajectories [17, 18] or training reinforcement learning (RL) policies for trajectory generation [19]. While BC is fast and straightforward to train, it cannot directly learn to minimize the optimization time of MPC. Instead, all supervised learning can do is to imitate the output of the solver. Worse yet, BC suffers from covariance shift, leading to poor generalization in zero-shot scenarios [20]. In contrast, RL is capable of directly optimizing performance objectives but requires extensive interaction data and struggles to learn from scratch [21]. As such, there remains a need to bridge these two paradigms—combining the efficiency of BC with the adaptability of RL—to effectively warm-start MPC across diverse solver families.
To overcome the limitations of prior work, we propose a learning framework that trains a policy to generate high-quality initial guesses for optimization-based control algorithms (e.g., MPC). The framework combines offline behavior cloning (BC) with online reinforcement learning (RL) fine-tuning and an additional zero-shot testing phase, as depicted in Fig. 1. The two training phases work synergistically. The online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimizing the optimization time of MPC. At the same time, the offline BC provides the online learning algorithm with a good starting point to expedite the training process. Moreover, initializing MPC with a learned policy rather than replacing MPC with an end-to-end controller preserves the original control formulation. The policy only sets the starting point, while MPC enforces system constraints to ensure safer operation. We demonstrate the generality of our approach by evaluating it across both deterministic MPC solver (COBYLA [22]) and sampling-based MPC solver (MPPI [23]). Our key contributions are:
  1. 1.
    We propose a novel two-phase learning framework to warm-start optimization-based control algorithms (e.g., MPC), enhancing both efficiency and robustness.
  2. 2.
    We empirically evaluate the framework on high-speed vehicle control tasks ( deterministic MPC), achieving $21.6\%$ faster optimization and $34.1\%$ higher tracking accuracy on zero-shot Formula 1 tracks compared to single-phase training.
  3. 3.
    We further test the framework on path-planning tasks (sampling-based MPC), showing $100\%$ improved safety, $12.8\%$ higher path efficiency, and $7.2\%$ smoother steering on zero-shot tracks compared to single-phase training.
II Preliminaries - Model Predictive Control
MPC is formulated as a finite-horizon optimization problem that minimizes a cost function while satisfying system dynamics and constraints. The MPC control law is formulated in Eqs.(1). At each time step $t$, MPC observes the current state $x_{t}$ and computes an optimal sequence of control inputs $U=[u_{t},\ldots,u_{t+H-1}]$ over a planning horizon $H$ by minimizing a cost function $J$ (Eq. (1a)). The system dynamics $x_{t+i+1}=f(x_{t+i},u_{t+i})$ (Eq. (1b)) predict how control inputs influence future states. Eq. (1c) and Eq. (1d) specify the constraints for the control inputs and states. After solving the optimization, only the first control input $u_{t}$ is applied, and the process repeats at the next time step using the updated state.
$J=\underset{U}{\text{minimize}}\sum_{i=0}^{H-1}l(x_{t+i},u_{t+i})$ (1a)
$\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
$U=[u_{t},...,u_{t+H-1}]\in\mathcal{U}_{j}\ \text{for all}\ j=1,\dots,n_{c_{u}}$ (1c)
$X=[x_{t},...,x_{t+H}]\in\mathcal{X}_{j}\ \text{for all}\ j=1,\dots,n_{c_{x}}$ (1d)
III Method
In this section, we discuss in detail the proposed two-phase training framework to learn a warm-start policy that reduces MPC solver runtime while maintaining tracking performance.. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
Algorithm 1 Offline Training
1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect NN
2:Initialize neural network policy $\pi_{\theta}^{\text{warm}}$
[…]
4:while $t<N$ do
5:  $(u_{t},u_{t+1},\cdots,u_{t+H-1})\leftarrow\pi^{\text{MPC}}(s,\vec{0})$
6:  $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s,u_{t},u_{t+1},\cdots,u_{t+H-1})\}$
7:  $s\leftarrow T(s,u_{t})$
8:  $t\leftarrow t+1$
9:end while
10:Train $\pi_{\theta}^{warm}$ with Eq. (2) and $\mathcal{D}$
III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs. The expert MPC, $\pi^{MPC}$, controls the agent to complete the task without a warm-started initial guess. At each step, $\pi^{MPC}$ observes a state $s$, takes an all-zero vector as the initial guess, and optimizes it to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). The state-action pair $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in $\mathcal{D}$ (line 6). The first action is then applied, transitioning the system to the next state based on environment transition $T$ (line 7–8). During data collection, $\pi^{MPC}$ runs enough iterations to ensure high control performance by disabling the “early stop.”
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [24]. Given the current vehicle state, the $\pi_{\theta}^{warm}$ predicts a sequence of actions that serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10), where BC learns a control policy $\pi_{\theta}$ from pre-collected MPC demonstrations by minimizing the mean squared error (MSE) between predicted and demonstrated actions, as shown in Eq. (2).
$\theta^{*}=\min_{\theta}{\sum_{\tau^{i}\in\mathcal{D}}\sum_{t=0}^{T}{(\pi_{\theta}^{warm}(s^{i})_{t}-u_{t}^{i})^{2}}}$ (2)
Algorithm 2 Online Fine Tuning
1:Input: fast MPC $\pi^{\text{MPC}}_{\text{fast}}$ with planning horizon $H$ and maximum optimization iteration $N_{fast}$, environment transition $T$, pre-trained warm-start policy $\pi_{\theta}^{warm}$
2:for each RL training iteration do
3:  Perceive an observation $s$
4:  $\hat{U}_{t}=(\hat{u}_{t},\hat{u}_{t+1},...,\hat{u}_{t+H-1})\leftarrow\pi^{warm}(s)$
5:  $pos_{i+1}^{car}=M_{dynamics}((s_{i},\hat{u}_{i}))|_{i=t}^{H-1}$
6:  Evaluate the quality of the planned trajectory
7:  $U_{t}^{\text{MPC}}=(u_{t},u_{t+1},...,u_{t+H-1})\leftarrow\pi^{\text{MPC}}_{\text{fast}}(s,\hat{U}_{t})$
8:  Calculate reward in RL using Eq. (8)
9:  $L_{imitation}=MSE(U_{t}^{MPC},U_{t}^{guess})$
10:  Compute training loss $L$ in Eq. (3)
11:  $s\leftarrow T(s,u_{t})$
12:  Update $\pi_{\theta}^{warm}$ with $L$
13:end for
III-B Online Fine Tuning
Our second phase, online fine-tuning, addresses the shortcomings of BC by combining the strengths of RL and Dataset Aggregation (DAgger) [25]. RL enables us to directly optimize the MPC computation time. Meanwhile, DAgger mitigates the distribution shift problem by augmenting the training set with additional expert demonstrations.
DAgger builds upon BC by incorporating online interaction with the environment and online querying of the expert. Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training. This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow A$ , is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
DAgger is integrated as a term in actor loss in the RL training, as shown in Eq. (3). $B$ is the batch size and $Q_{\theta}$ is the critic network. $L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. DAgger loss is only added when the quality of the initial guess is lower than a certain threshold. To reduce the overhead of repeatedly querying the expert MPC, $L_{imitation}$ is only applied when the initial guess falls below a certain threshold. This accelerates RL training while maintaining regularization from expert demonstrations. $\lambda$ is the weight coefficient between the RL loss and imitation loss.
$L$ $=\lambda\cdot L_{actor}+(1-\lambda)\cdot L_{imitation}$ (3)
$L_{\text{actor}}$ $=-\frac{1}{B}\sum_{i=1}^{B}Q_{\theta}(s_{i},\pi^{\text{warm}}(s_{i}))\cdot L_{value}$
IV EXPERIMENTS SETUP
In this section, we discuss in detail the experimental setup for both deterministic and sampling-based MPC. For deterministic MPC, we evaluate the path-tracking problem on high-speed Formula 1 tracks [26], where minimizing optimization time is critical for maintaining performance. For sampling-based MPC, we choose the navigation problem in environments without a pre-defined reference path, requiring the vehicle to generate its own collision-free trajectory.
IV-A Deterministic MPC
IV-A1 MPC Details
The Formula 1 domain consists of three training and seven zero-shot testing tracks, as shown in Fig. 2(a) and Fig. 2(b). The training tracks are used for demonstration collection and policy learning, while the zero-shot tracks remain unseen during training to test generalization. Each track’s reference trajectory is defined by waypoints along the centerline and scaled down 10:1 to make each lap a reasonable length. The friction between the tire and the road is not considered in our dynamics model.
(a) Training tracks
(b) Zero-shot tracks
Figure 2: Training and Testing maps used in the Formula 1 track path-tracking domain.
$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt+\epsilon_{1}$ (4a)
$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt+\epsilon_{2}$ (4b)
$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt+\epsilon_{3}$ (4c)
$v_{t+1}=\begin{cases}v_{t}+a_{t}\,\Delta t+\epsilon_{4},&\text{if }v_{t}+a_{t}\,\Delta t<10,\\ 10,&\text{otherwise.}\end{cases}$ (4d)
$\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I),\quad\boldsymbol{\epsilon}=[\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}]^{\top}$ (4e)
MPC controls vehicle acceleration and steering to follow the reference trajectory using the dynamics model $M_{dynamics}$ in Eq. (4). $(x_{t}^{car},y_{t}^{car})$ denote the vehicle’s global position; $v_{t}$ and $yaw_{t}$ are its speed ($m/s$) and yaw angle (radrad); $a_{t}$ and $\theta_{t}^{\text{steering}}$ represent acceleration ($m/s^{2}$) and steering angle (radrad). The wheelbase is L=2.89mL=2.89~m, and the time step is dt=0.02sdt=0.02~s. Gaussian noise $\epsilon_{1}$ - $\epsilon_{4}$ with $\sigma=0.01$ is added to model uncertainty.
The MPC objective function is composed of five parts as shown in Eq. (5). The first two terms are Cross Track Error (xtexte) and Error in Heading (etheth) computed using Eq. (6). and Eq. (7) respectively. etheth denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. $v_{t}^{\text{ref}}=10m/s$ is the desired speed of the vehicle. The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds.
$J_{t}$ $=\sum_{i=t}^{t+H-1}(w_{0}\cdot xte_{i}^{2}+w_{1}\cdot eth_{i}^{2}$ (5)
$\quad+w_{2}\cdot(v_{i}-v_{t}^{\text{ref}})^{2}$
$\quad+w_{3}\cdot(steer_{i}-steer_{i-1})^{2}$
$\quad+w_{4}\cdot(throttle_{i}-throttle_{i-1})^{2})$
$xte=distance(pos^{car}_{t},\mathbf{wp}_{i}(x_{i},y_{i})^{closest})$ (6)
$eth=abs\left(yaw^{car}_{t}-\arctan\left({\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}}}\right)\right)$ (7)
Gradient-based solvers are highly effective when both system dynamics and cost functions are smooth and differentiable. However, in our setting, the calculation of xtexte (cross-track error) requires finding the closest point on a reference trajectory to the vehicle’s current position, which introduces discontinuities and non-differentiability into the objective function. This motivates the use of gradient-free solvers such as COBYLA [22], which are better suited for non-differentiable optimization problems [27, 28]. Nonetheless, with high speeds and on tracks with sharp turns, gradient-free MPC solvers struggle to optimize trajectories in real time without a good initial guess, making practical deployment challenging. As shown in Section V, the gradient-free MPC solver succeeds only on the simple IMS track but fails on all complex tracks.
We implement an early-stop condition when the planned trajectory’s accumulated xtexte is below $0.1,m$, preventing MPC wasting time on optimizing a trajectory that is already good enough. We also cap MPC iterations at 100 during RL training to avoid solver stagnation and evaluate performance under both 100 and 200 iteration limits during testing for sensitivity analysis.
IV-A2 RL Details
The reward at each step of the RL training is shown in Eq. (8). The first term is the negative MPC optimization time, and the second term is the negative of the xtexte over the planning horizon, $H$. This reward design helps optimize $\pi_{\theta}^{warm}$ by minimizing the MPC running time directly while maximizing tracking accuracy.
$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (8)
IV-B Sampling Based MPC
IV-B1 MPC Details
Unlike deterministic MPC, MPPI employs a sampling-based stochastic optimization framework to minimize a predefined cost. This setup allows us to test whether our warm-start algorithm can also enhance a different control optimizer. As MPPI is mainly used for path-planning and obstacle-avoidance tasks [29, 30], we build an obstacle-rich navigation domain for evaluation, as shown in Fig. 3. The training set includes two straight and two curved-road maps, while the testing set contains three unseen vertical and three unseen curved-road maps. Each map features randomly generated polygonal obstacles with varying sizes, shapes, and positions, posing diverse planning challenges. The vehicle starts at a green dot and must reach the goal region (red dot) without collisions. To increase training diversity, each training map offers two initialization points: one on the left and one on the right.
$J=\sum_{i=t}^{t+H-1}$ $\Big[w_{1}\,throttle_{i}^{2}+w_{2}\,steer_{i}^{2}$ (9)
$\quad+w_{3}\,(v_{i}-v_{t}^{ref})^{2}$
$\quad+w_{4}\,C_{obstacle}(x_{i},y_{i})$
$\quad+w_{5}\,C_{border}(x_{i},y_{i})\Big]$
$+\quad\sum_{i=t+H-3}^{t+H-1}\Big[w_{6}\,C_{goal}(x_{i},y_{i})\Big]$
The vehicle dynamics model is identical to that employed in the deterministic MPC in Eq. (4), ensuring a fair comparison. The cost function for MPPI is shown in Eq. (9). The first two terms correspond to smoothness penalties to enhance control stability. $C_{obstacle}$ penalizes proximity to obstacles, $C_{border}$ penalizes deviations beyond road boundaries to maintain lane centering, and $C_{goal}$ penalizes the final distance to the target region. The planning horizon is fixed at 25 steps. And number of samples is set to 200 samples to balance efficiency and control quality.
(a) Training Maps
(b) Testing Maps
Figure 3: Training and testing maps in the obstacle-rich navigation domain. Green dots indicate spawn points, red dots represent destinations, and polygons correspond to obstacles.
IV-B2 RL Details
The RL reward function is defined in Eq. (10). It is aligned with the MPPI objective function. The only difference is the final term, which encourages the vehicle to reduce its distance to the goal at each timestep. Since MPPI operates via parallel trajectory sampling rather than iterative optimization, we exclude the MPC optimization time from the RL reward, as all trajectories are sampled simultaneously on the GPU. However, warm-starting remains valuable for onboard systems with limited parallelism. A good initial guess allows the controller to achieve comparable performance with fewer samples or higher accuracy with the same number of samples. In our evaluation, we fix the number of samples and test whether our method outperforms the baselines in control performance.
$r_{t}^{mppi}$ $=-w_{1}\,throttle_{t}^{2}-w_{2}\,steer_{t}^{2}$ (10)
$\quad-w_{3}\,(v_{t}-v_{t}^{ref})^{2}-w_{4}\,C_{obstacle}(x_{t},y_{t})$
$\quad-w_{5}\,C_{border}(x_{t},y_{t})$
$\quad-w_{6}\,((C_{goal}(x_{t},y_{t})-(C_{goal}(x_{t-1},y_{t-1}))$
TABLE I: Deterministic MPC results with a maximum of 100 optimization iterations. Values represent averages over three test runs with different random seeds. Reported percentages indicate relative improvement compared to the corresponding baseline. The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s. means no significance.
All Zeros &
Previous Solution &
Ours w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Training Tracks
Zandvoort 0.1270 ± 0.0800 0.5929 ± 0.6066 0.1530 ± 0.0542 0.3869 ± 0.2840 0.1087 ± 0.0821 0.3975 ± 0.3084
Sochi 0.1429 ± 0.0999 0.5657 ± 0.5869 0.1772 ± 0.0620 0.4704 ± 0.5150 0.1128 ± 0.1010 0.3755 ± 0.3075
Hockenheim Failed 0.1129 ± 0.0912 0.4686 ± 0.4177 0.1450 ± 0.0581 0.4078 ± 0.4665 0.0865 ± 0.0877 0.3548 ± 0.3373
Training Tracks
Total
0.1285 ± 0.0919 0.5450 ± 0.5504 0.1597 ± 0.0600 0.4246 ± 0.4385 0.1033 ± 0.0920 0.3758 ± 0.3179
Compared with
Ours w/o RL
-24.27%
n.s.
22.10%
19.65%
30.04%
Zero-shot Tracks
Nuerburgring 0.1845 ± 0.0669 1.6630 ± 1.5488 0.1679 ± 0.0601 0.3774 ±0.3233 0.1158 ± 0.0967 0.4090 ± 0.3500
BrandsHatch 0.1163 ± 0.0922 0.5117 ± 0.5484 0.1545 ± 0.0519 0.3463 ± 0.2075 0.0895 ± 0.0837 0.3355 ± 0.2592
Oschersleben 0.1195 ± 0.0832 0.4181 ± 0.3514 0.1421 ± 0.0582 0.4383 ± 0.4309 0.1116 ± 0.0846 0.3405 ± 0.2395
Shanghai Failed 0.1845 ± 0.0732 1.3278 ± 1.2934 0.1717 ± 0.0602 0.4242 ± 0.4157 0.1189 ± 0.0977 0.3869 ± 0.2688
Melbourne 0.1375 ± 0.0963 0.7781 ± 0.9978 0.1767 ± 0.0588 0.4297 ± 0.3378 0.1235 ± 0.0937 0.3921 ± 0.3028
Catalunya 0.1140 ± 0.0942 0.5561 ± 0.6710 0.1561 ± 0.0571 0.4095 ± 0.3798 0.1019 ± 0.0891 0.3631 ± 0.2863
Zero-shot Tracks
Total
0.1558 ± 0.0871 1.0696 ± 1.2447 0.1629 ± 0.0589 0.4033 ± 0.3539 0.1095 ± 0.0913 0.3707 ± 0.2905
Compared with
Ours w/o RL
-4.57%
n.s.
62.30%
29.75%
65.34%
TABLE II: Deterministic MPC results with a maximum of 200 optimization iterations. Values represent averages over three test runs with different random seeds. Reported percentages indicate relative improvement compared to the corresponding baseline. The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s. means no significance.
All Zeros &
Previous Solution &
Ours w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Optimization
Time (Second)
xte
Training Tracks
Zandvoort 0.2395 ± 0.1700 0.5130 ± 0.5054 0.2498 ± 0.1412 0.3472 ± 0.2664 0.1954 ± 0.1753 0.3501 ± 0.2402
Sochi 0.2821 ± 0.2079 0.5367 ± 0.5056 0.2836 ± 0.1649 0.3783 ± 0.3697 0.2166 ± 0.2122 0.3621 ± 0.2745
Hockenheim Failed 0.2072 ± 0.1914 0.4245 ± 0.3663 0.2349 ± 0.1502 0.3754 ± 0.4123 0.1593 ± 0.1815 0.3172 ± 0.2734
Training Tracks
Total
0.2451 ± 0.1935 0.4944 ± 0.4695 0.2578 ± 0.1545 0.3674 ± 0.3546 0.1918 ± 0.1930 0.3441 ± 0.2645
Compared with
Ours w/o RL
-5.19%
n.s.
25.70%
21.77%
30.41%
Zero-shot Tracks
Nuerburgring 0.3086 ± 0.1865 0.5814 ± 0.5536 0.2685 ± 0.1576 0.3359 ±0.3251 0.2358 ± 0.2008 0.4254 ± 0.3867
BrandsHatch 0.2172 ± 0.1887 0.4122 ± 0.3194 0.2514 ± 0.1421 0.3063 ± 0.1850 0.1584 ± 0.1771 0.3213 ± 0.2282
Oschersleben 0.2340 ± 0.1754 0.3977 ± 0.3196 0.2467 ± 0.1473 0.3623 ± 0.2838 0.2376 ± 0.1756 0.3893 ± 0.2840
Shanghai Failed 0.3093 ± 0.1943 0.6666 ± 0.6522 0.2806 ± 0.1551 0.3402 ± 0.3107 0.2150 ± 0.2028 0.3765 ± 0.2549
Melbourne 0.2747 ± 0.2011 0.6036 ± 0.6359 0.2930 ± 0.1561 0.3868 ± 0.3315 0.2223 ± 0.2016 0.3584 ± 0.2459
Catalunya 0.2300 ± 0.1947 0.5820 ± 0.6688 0.2663 ± 0.1521 0.3896 ± 0.3693 0.1920 ± 0.1871 0.3704 ± 0.3305
Zero-shot Tracks
Total
0.2663 ± 0.1949 0.5549 ± 0.5710 0.2697 ± 0.1531 0.3543 ± 0.3112 0.2078 ± 0.1921 0.3689 ± 0.2909
Compared with
Ours w/o RL
-1.30%
n.s.
36.16%
21.97%
33.52%
IV-C Baselines and Metrics
For deterministic MPC, we evaluate performance using two metrics: (1) the average optimization time per step (seconds) and (2) the average xtexte (meters) per step. For sampling-based MPC, we employ three metrics: (1) steering standard deviation, to assess control stability; (2) average number of steps required to reach the goal, as a proxy for path efficiency; and (3) number of obstacle collisions, to evaluate safety and avoidance capability. During testing, we compare the performance of our algorithm against five types of initial-guess policies:
  • All Zero: all-zero initial guesses.
  • Previous Solution: initial guesses derived from the MPC solution at the previous step.
  • Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training.
  • Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC.
  • Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger.
To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
TABLE III: Sampling Based MPC Results. Values represent averages over three test runs with different random seeds. Numbers in parentheses denote the standard error.
All Zeros Previous Solution
Ours
w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Map
Name
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Training Maps
Vertical_1 0.6647
287.3333
± 22.8356
2.0000
± 1.4142
0.6306
251.8333
± 12.0568
0.0000
± 0.0000
0.4881
258.0000
± 19.7281
1.0000
± 1.0954
0.4747
278.1667
± 16.7978
0.6667
± 0.8165
0.4712
245.1667
± 20.8750
0.8333
± 1.6021
0.4689
249.3333
± 22.2860
0.0000
± 0.0000
Vertical_2 0.6491
289.0000
± 45.4093
0.6667
± 1.2111
0.6393
273.6667
± 11.5873
0.0000
± 0.0000
0.4923
264.0000
± 23.0304
0.8333
± 0.9832
0.4637
279.8333
± 24.7258
0.0000
± 0.0000
0.4810
265.6667
± 19.2942
0.0000
± 0.0000
0.4716
256.5000
± 26.8384
0.1667
± 0.4082
Curved_1 0.6290
369.8333
± 37.0751
1.5000
± 3.6742
0.6009
360.3333
± 13.4263
0.0000
± 0.0000
0.4773
376.3333
± 84.5167
0.0000
± 0.0000
0.4748
354.6667
± 36.9035
0.0000
± 0.0000
0.4322
334.6667
± 4.1793
0.0000
± 0.0000
0.4256
335.0000
± 5.6569
0.0000
± 0.0000
Curved_2 0.6366
422.6667
± 65.9019
9.3333
± 17.5803
0.6369
401.3333
± 26.1738
0.0000
± 0.0000
0.5064
427.6667
± 60.8857
1.0000
± 2.4495
0.4879
460.5000
± 81.4389
0.0000
± 0.0000
0.4812
375.6667
± 24.7117
1.3333
± 2.0656
0.4800
376.1667
± 17.6682
0.0000
± 0.0000
Training Map
Total
0.6433
342.2083
± 72.2255
3.3750
± 9.1358
0.6266
321.7917
± 64.5587
0.0000
± 0.0000
0.4922
331.5000
± 89.9633
0.7083
± 1.3981
0.4775
343.2917
± 87.7870
0.1667
± 0.4815
0.4663
305.2917
± 56.4088
0.5417
± 1.3507
0.4619
304.2500
± 57.5917
0.0417
± 0.2041
Testing Maps
Vertical_1 0.6425
285.6667
± 107.7698
9.6667
± 7.7675
0.6228
288.3333
± 97.6183
0.0000
± 0.0000
0.4667
225.0000
± 1.0000
0.0000
± 0.0000
0.4982
355.6667
± 198.4574
0.0000
± 0.0000
0.4697
230.0000
± 21.7945
0.3333
± 0.5774
0.4579
220.6667
± 4.5093
0.0000
± 0.0000
Vertical_2 0.6416
266.3333
± 35.5575
2.0000
± 2.6458
0.6373
231.6667
± 3.0551
0.0000
± 0.0000
0.5068
235.6667
± 14.2244
1.0000
± 1.0000
0.4709
225.3333
± 4.1633
0.6667
± 1.1547
0.4592
217.6667
± 3.2146
0.0000
± 0.0000
0.4421
219.3333
± 14.4338
0.0000
± 0.0000
Vertical_3 0.6298
312.0000
± 106.2262
3.3333
± 4.9329
0.6280
263.6667
± 38.5011
0.0000
± 0.0000
0.4332
220.3333
± 1.1547
0.0000
± 0.0000
0.4646
258.0000
± 4.3589
0.0000
± 0.0000
0.4685
257.3333
± 1.5275
0.0000
± 0.0000
0.4621
239.0000
± 25.1197
0.0000
± 0.0000
Curved_1 0.6251
348.0000
± 2.6458
9.6667
± 12.6623
0.5929
328.6667
± 3.0551
0.0000
± 0.0000
0.4457
323.6667
± 3.5119
0.0000
± 0.0000
0.4385
328.0000
± 10.5357
0.0000
± 0.0000
0.4267
317.6667
± 0.5774
0.0000
± 0.0000
0.4227
315.6667
± 1.1547
0.0000
± 0.0000
Curved_2 0.5890
331.0000
± 9.1652
0.3333
± 0.5774
0.5720
319.6667
± 0.5774
0.0000
± 0.0000
0.4163
314.0000
± 1.7321
0.0000
± 0.0000
0.4060
310.6667
± 1.1547
0.0000
± 0.0000
0.4130
308.6667
± 2.0817
0.0000
± 0.0000
0.4120
310.0000
± 4.5826
0.0000
± 0.0000
Curved_3 0.6416
418.0000
± 83.5165
3.3333
± 3.5119
0.5841
339.3333
± 8.1445
0.0000
± 0.0000
0.4654
340.6667
± 26.3881
0.0000
± 0.0000
0.4975
387.6667
± 69.8164
1.3333
± 2.3094
0.4446
324.3333
± 5.6862
0.0000
± 0.0000
0.4330
326.0000
± 7.2111
0.0000
± 0.0000
Testing Map
Total
0.6284
326.8333
± 78.7821
4.7222
± 6.7196
0.6050
295.2222
± 53.3728
0.0000
± 0.0000
0.4550
276.5556
± 52.8597
0.1667
± 0.5145
0.4713
310.8889
± 91.9545
0.3333
± 1.0290
0.4467
275.9444
± 44.7667
0.0556
± 0.2357
0.4372
271.7778
± 48.6047
0.0000
± 0.0000
TABLE IV: Relative improvement between different methods in sampling-based MPC and statistical test results. The reported percentages indicate relative improvement. The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s. means no significance.
Previous Solution
Ours
w/o BC
Ours
w/o RL
Ours
w/o DAgger
Ours
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Steer
Std
Number
of Steps
Number
of Hits
Training Maps
Compared with
All Zeros
2.60%
5.97%
100%
23.49%
3.13%
n.s.
79.01%
25.77%
-0.32%
n.s.
95.06%
27.51%
10.79%
83.95%
28.20%
11.09%
98.76%
Compared with
Previous Solution
21.45%
-3.02%
n.s.
23.80%
-6.68%
n.s.
25.58%
5.13%
26.27%
5.46%
Compared with
Ours w/o BC
2.99%
-3.56%
n.s.
76.46%
5.26%
7.91%
23.52%
n.s.
6.16%
8.22%
94.11%
Compared with
Ours w/o RL
2.35%
n.s.
11.07%
-224.96%
n.s.
3.27%
n.s.
11.37%
74.99%
n.s.
Compared with
Ours w/o DAgger
0.95%
n.s.
0.34%
n.s.
92.30%
Testing Maps
Compared with
All Zeros
3.72%
9.67%
100.00%
27.59%
15.38%
96.47%
25.00%
4.88%
92.94%
28.91%
15.57%
98.82%
30.43%
16.85%
100.00%
Compared with
Previous Solution
24.79%
6.32%
22.10%
-5.31%
n.s.
26.17%
6.53%
27.74%
7.94%
0.00%
Compared with
Ours w/o BC
-3.58%
n.s.
-12.41%
n.s.
-99.94%
n.s.
1.82%
n.s.
0.22%
n.s.
66.65%
n.s.
3.91%
1.73%
100.00%
n.s.
Compared with
Ours w/o RL
5.21%
n.s.
11.24%
83.32%
n.s.
7.22%
12.58%
100.00%
n.s.
Compared with
Ours w/o DAgger
2.13%
1.51%
n.s.
100.00%
n.s.
V Experiment Results
In this section we design experiments to test the following research questions in both deterministic MPC and sampling-based MPC problem:
  • RQ1 Does our warm-start policy reduce solver runtime in deterministic MPC problem?
  • RQ2 Does our warm-start policy improve tracking performance in deterministic MPC problem?
  • RQ3 Does our warm-start policy generalize to unseen tracks in deterministic MPC problem?
  • RQ4 Does our warm-start policy generalize from deterministic MPC to sampling-based MPC problem?
V-A Deterministic MPC
We perform testing on both the training tracks (Fig. 2(a)) and the challenging zero-shot tracks (Fig. 2(b)). The results are shown in Table I. On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy. Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely via offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality and covariance shift problem in BC.
On the training tracks, compared to Ours w/o RL, the proposed method improves optimization time by $19.95\%$ and tracking accuracy by $30.60\%$. When compared to Ours w/o RL, Ours w/o DAgger achieves a $19.12\%$ improvement in tracking accuracy; however, the optimization time degrades by $22.08\%$, highlighting the necessity of incorporating the DAgger training loss during RL training.
On the zero-shot tracks, compared to Ours w/o RL, the proposed method improves optimization time by $21.63\%$ and tracking accuracy by $34.12\%$. Similarly, Ours w/o DAgger achieves a $30.35\%$ improvement in tracking accuracy, but optimization time degrades by $18.00\%$. This degradation pattern aligns with that on the training tracks, further highlighting the importance of integrating the DAgger training loss during RL training.
Figure 4: Visualization of optimized trajectories for our method and the baselines on a segment of the Nürburgring map.
The results also reveal that initializing an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane when facing any sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 100 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
We also relaxed the maximum MPC iterations to 200 and repeated the test. The results in Table II show that our warm-start policy again achieves the greatest improvement in optimization time and the largest performance gains over the baselines, further supporting RQ1, RQ2, and RQ3.
A visualization of the optimized trajectories for our method and the baselines on a segment of the Nürburgring map is presented in Fig. 4 (MPC optimization iterations=100). The figure illustrates that using our method as a warm-start strategy enables the MPC to track the reference path more accurately by tracking the sharp more accurately.
V-B Sampling-Based MPC
We evaluate performance on the training maps (Fig. 3(a)) and the zero-shot maps (Fig. 3(b)), with results summarized in Table III. Consistent with the findings for deterministic MPC, our proposed framework delivers the most significant performance improvements for the MPPI controller compared to all baselines. Notably, the improvements are more pronounced in the zero-shot environments, demonstrating the strong generalizability of the proposed method.
V-C Findings
In summary, our empirical results support that:
  • Our proposed framework, trained offline or with online fine-tuning, outperform other non-learning-based baselines.
  • The two-phase training algorithm generalizes better on zero-shot tracks, achieving superior zero-shot performance.
  • Offline training effectively accelerates online fine-tuning, reducing training time to achieve strong performance.
VI LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, it also has several limitations for future research that merit consideration. First, our experiments assume access to perfect state information. Second, the current evaluation is limited to vehicle control. Future work could expand into higher-DOF robotic platforms such as manipulators or aerial robots.
VII CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver. Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time without degrading MPC’s performance. This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.